OpenStudy (anonymous):

Reciprocal of 2cos(2x)? Help for a medal?

OpenStudy (mathmale):

Skip the medal part; it's unnecessary. What does reciprocal mean to you? What is the reciprocal of 3?

OpenStudy (anonymous):

The reciprocal of 3 is 1/3 I believe

OpenStudy (mathmale):

right. What is the reciprocal of cos x?

OpenStudy (anonymous):

1/secx

OpenStudy (mathmale):

why do you believe division is necessary here?

OpenStudy (mathmale):

"reciprocal of cos x" looks like 1 / cos x

OpenStudy (anonymous):

I just have the reciprocal identities memorized

OpenStudy (anonymous):

at least I thought I did

OpenStudy (mathmale):

The reciprocal of cos x is .... what? no fraction necessary.

OpenStudy (mathmale):

Either memorize the fact that the rec. of cos x is sec x, or actually write 1 / cos x and recall that the equivalent of 1 / cos x is sec x.

OpenStudy (mathmale):

The reciprocal of sin x is 1 / sin x, or csc x.

OpenStudy (mathmale):

The reciprocal of tan x is 1 / tan x, or cot x.

OpenStudy (mathmale):

The recip of cos x is .... ?

OpenStudy (anonymous):

1/cosx

OpenStudy (mathmale):

Yes, and that is also equal to ... ????

OpenStudy (anonymous):

secx

OpenStudy (mathmale):

Very good!

OpenStudy (anonymous):

I've used this reference in the past (referring to the 2nd page). http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf Is it not correct?

OpenStudy (mathmale):

Reciprocal of 2cos(2x)?

OpenStudy (anonymous):

2sec2x?

OpenStudy (anonymous):

or 2(1/cos(2x))

OpenStudy (mathmale):

Your table of formulas is great, just fine, correct.

OpenStudy (anonymous):

Wait...so cosx reciprocal can also been represented as 1/secx?

OpenStudy (mathmale):

Reciprocal of 2cos(2x)? You have to put all of 2cos(2x) in the denominator. 1 / ( ? )

OpenStudy (mathmale):

cosx = 1 / sec x, YES sec x = 1 / cos x YES

OpenStudy (mathmale):

Reciprocal of 2cos(2x)?

OpenStudy (anonymous):

1 / ( sec(2x) )

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I take that back

OpenStudy (mathmale):

OK

OpenStudy (mathmale):

try again

OpenStudy (anonymous):

2 / ( sec(2x) )

OpenStudy (mathmale):

Reciprocal of 2cos(2x)? All of " 2cos(2x)" goes into the DENOMINATOR.

OpenStudy (anonymous):

Oh so it's 1 / (2cos(2x)).

OpenStudy (mathmale):

Right. Or, (1/2)(1 / cos 2x )

OpenStudy (mathmale):

What is the trig identity for 1 / cos 2x?

OpenStudy (anonymous):

cos2x

OpenStudy (mathmale):

Sorry, I didn't think. Try answering my question again.

OpenStudy (mathmale):

Review our past conversation.

OpenStudy (mathmale):

cos 2x is not the reciprocal of cos 2x.

OpenStudy (mathmale):

Reciprocal of 2cos(2x)?

OpenStudy (anonymous):

"cos 2x is not the reciprocal of cos 2x" That I understand. "Try answering my question again." Which question?

OpenStudy (mathmale):

Reciprocal of 2cos(2x)?

OpenStudy (anonymous):

1 / (2cos(2x)) or 1 / (2sec(2x)) or 2sec(2x)

OpenStudy (anonymous):

correct?

OpenStudy (mathmale):

Actually, (1/2)sec (2x). This follows the pattern we established earlier.

OpenStudy (mathmale):

recall that the recip. of 2 is (1/2).

OpenStudy (anonymous):

Apologies. My page keeps reloading itself, and I'm not sure why.

OpenStudy (mathmale):

Same here.

OpenStudy (mathmale):

Reciprocal of 2cos(2x)? Find the recip. of 2 and then the recip of cos 2x, and mult. the 2 recips together.

OpenStudy (anonymous):

So it can also be (1/2)(1/cos(2x)) = 1/(2cos(2x))

OpenStudy (mathmale):

Your first form is correct, and so is the second. However, it's better to write (1/2)sec 2x.

OpenStudy (anonymous):

So for the reciprocals of cosx... 1/cos(x) and sec(x) both work, but 1/sec(x) is the best way to show it?

OpenStudy (mathmale):

1/sec(x) is the best way to show it? No. recall that 1/cos x = sec x.

OpenStudy (mathmale):

You were finding the recip of 2 cos 2x, which is (1/2)*(1/cos 2x) = (1/2)(sec 2x) END

OpenStudy (anonymous):

Yes. I was just using the cos(x) as an example.

OpenStudy (mathmale):

Thx for your patience. Clear on this? Further questions?

OpenStudy (anonymous):

I think I got it. Thank you!

OpenStudy (mathmale):

You're very welcome. Take care!

OpenStudy (mathmale):

Thx for the3 medal!

OpenStudy (anonymous):

I know you said no medal, but I'm thankful. You're welcome. :3

OpenStudy (mathmale):

See y ou again! Bye.

OpenStudy (anonymous):

Good bye. :)